OpenStudy (anonymous):

HELP ME PLEASE!!! Write an explicit formula for the following sequence. 1/2, 1/4, 1/6, 1/8, 1/10. . . Find the twentieth term.

OpenStudy (anonymous):

the \(\color{red}{1}\)st term: \(\dfrac{1}{2*\color{red}{1}}\) the \(\color{blue}{2}\)nd term: \(\dfrac{1}{2*\color{blue}{2}}\) the \(\color{green}{3}\)rd term: \(\dfrac{1}{2*\color{green}{3}}\) then \(\color{red}{4}\)st term: \(\dfrac{1}{2*\color{red}{4}}\) can you write down the 20th term?

OpenStudy (anonymous):

*4th

OpenStudy (anonymous):

So it would be 1/40? But how do I put it in explicit form?

OpenStudy (anonymous):

Like the formula is \[a_{n}=a_{1}+d(n-1)\] I just do not know where to put the numbers asub1 would be the first sequence, d is the common difference and n is 20

OpenStudy (anonymous):

just replace the number by letter n that is \(a_n= \dfrac{1}{2n}\)

OpenStudy (anonymous):

Ohhh okay i understand now!! I just had to look for the twentieth term and not the equation for it. I just got confused. THANK YOU SO MUCH!!